What is the specific heat of aluminum if the temperature of a 28.4 g sample of aluminum is increased by 8.1 C when 207j of heat is added
1
Expert's answer
2017-05-18T06:21:09-0400
Q = cmΔT
m (Al) = 28.4/1000 = 0.0284 kg
c (Al) = Q/(mΔT) = 207/(0.0284·8.1) = 900 J/(kg⋅K)